# Variational Stochastic Gradient Descent for Deep Neural Networks

### v1

@@ 4 Experiments
Figure 2: Top-1 test accuracy on CIFAR100 dataset.
Figure 2: Top-1 test accuracy on CIFAR100 dataset.
Figure 3: Top-1 test accuracy on TinyImagenet-200 dataset.
In this Section, we evaluate the performance of the VSGD compared to other optimizers on high-dimensional optimization problems  [^1] Code is available at
https://github.com/generativeai-tue/vsgd
. The task we focus on in our work is training DNNs to solve classification tasks.
4.1 Setup
Data
We used three benchmark datasets: CIFAR100 (Krizhevsky et al., 2009), TinyImagenet-200 (Deng et al., 2009b) [^2] We use Tiny Imagenet(Stanford CS231N) provided on www.image-net.org, and Imagenet-1k (Deng et al., 2009a). The CIFAR100 dataset contains 60000 small ($32\times 32$) RGB images labeled into 100 different classes, 50000 images are used for training, and 10000 are left for testing. In the case of TinyImagenet-200, the models are trained on 100000 images from 200 different classes and tested on 10000 images. For a large-scale experiment, we use the Imagenet-1k dataset which contains 1,281,167 images from 1000 classes.
Architectures
Hyperparameters
4.2 Results
@@
ConvMixer-256/8 0.6 0.36 0.38
@@
In order to evaluate VSGD’s performance in heavily overparameterized domains, we performed an experiment on the Imagenet-1k dataset (Deng et al., 2009a) where we trained a ResNet-50 with both VSGD and Adam. Due to limited computational resources, we did not perform any hyperparameter tuning for this experiment. For both optimizers, we used a learning rate of 0.1 with a OneCycleLR learning rate scheduler (Smith & Topin, 2019) and a batch size of 256. As in the previous experiments, we set $\gamma$ to 1e-8. We train both models for 50 epochs. The top-1 accuracy of the validation set for VSGD was 71.1%, exceeding Adam’s 68.7% by 2.4%.
@@ 5 Conclusion
Broader Impact Statement
This paper presents work whose goal is to improve the field of optimization. As is the case with other optimization methods, the main potential societal risk of our work is the use of our approach for training deep neural network architectures that are to be used for harmful purposes.
@@ A.1 Derivation of VSGD
    Pick any $\phi_{t,k}=(\mu_{t,k},\sigma^{2}_{t,k})^{T}\in\phi_{t,1:T}$, and set $u_{t+1}=\mu_{t,k}$, where $u_{t+1}$ serves as the control variable for the next iteration. We can choose any $\phi_{t,k}$ because by the definition (Eq. 57), $\phi_{t,k}$ and $\phi_{t,j}$ are identical for any $i,j\in 1:T$.
@@ A.2 A Kernel Smoothing View on VSGD
There are two noisy observations of the gradient function at $\theta=\theta_{t}$. They are $(\theta_{t},u_{t})$ and $(\theta_{t},\hat{g}_{t})$. We can define two Gaussian kernels for each of the noisy observations, governed by $w_{g}^{-1}$ and $w_{\hat{g}}^{-1}$. By denoting the Gaussian kernels $\kappa_{1}$ and $\kappa_{2}$, the smoothed gradient value at $\theta_{t}$ should be
$g_{t}=\frac{\kappa_{1}(\theta_{t},\theta_{t})}{\kappa_{1}(\theta_{t},\theta_{t})+\kappa_{2}(\theta_{t},\theta_{t})}u_{t}+\frac{\kappa_{2}(\theta_{t},\theta_{t})}{\kappa_{1}(\theta_{t},\theta_{t})+\kappa_{2}(\theta_{t},\theta_{t})}\hat{g}_{t}$ (70)
@@
Note that $u_{t}$ in VSGD is introduced as a learned summary of the noisy observations before $t$ and we only need to smooth on $u_{t}$ as a cheap approximation to smoothing on all of the historical data. Similar simplification technique is commonly seen in scalable Gaussian Process solutions, as discussed in (Quinonero-Candela & Rasmussen, 2005; Liu et al., 2020a). In these references, the learned summaries are called the ”inducing points”.
@@ A.3 Treating Samples Separately in a Mini-bach
Assuming the size of the mini-batch is $M$, denote the sample gradients $\{\hat{g}_{t}^{(i)}\}_{i=1:M}$. According to (Hoffman et al., 2013), we would need to calculate Eq. 15 $M$ times and average them in Eq. 18 and Eq. 19. Specifically, we would need to replace Eq. 15, Eq. 18 and Eq. 19 with:
@@ B.3 Data Augmentations
B.4 Training Curves
Figure 4: Training loss on CIFAR100 dataset.

### v2

@@ 4.2 Results
Figure 2: Top-1 test accuracy on CIFAR100 dataset.
Figure 3: Top-1 test accuracy on TinyImagenet-200 dataset.
@@
In order to evaluate VSGD’s performance in overparameterized domains, we performed an experiment on the Imagenet-1k dataset (Deng et al., 2009b) where we trained a ResNet-50 with both VSGD and Adam. Due to limited computational resources, we did not perform any hyperparameter tuning for this experiment. For both optimizers, we used a learning rate of 0.1 with a OneCycleLR learning rate scheduler (Smith & Topin, 2019) and a batch size of 256. Same as our previous experiments, we set $\gamma$ to 5e-8 for VSGD. Both models were trained for 50 epochs. The top-1 accuracy of the validation set for VSGD was 72.7%, exceeding Adam’s 69.8% by 2.9%. We also notice slight improvement in convergence speed (Figure 6).
@@ A.1 Derivation of VSGD
    Pick any $\phi_{t,k}=(\mu_{t,k},\sigma^{2}_{t,k})^{T}\in\phi_{t,1:T}$, and set $u_{t+1}=\mu_{t,k}$, where $u_{t+1}$ serves as the control variate for the next iteration. We can choose any $\phi_{t,k}$ because by the definition (Eq. 54), $\phi_{t,k}$ and $\phi_{t,j}$ are identical for any $i,j\in 1:T$.
@@ A.2 A Kernel Smoothing View on VSGD
There are two noisy observations of the gradient function at $\theta=\theta_{t}$. They are $(\theta_{t},u_{t})$ and $(\theta_{t},\hat{g}_{t})$. We can define two Gaussian kernels for each of the noisy observations, governed by $w_{g}^{-1}$ and $w_{\hat{g}}^{-1}$. By denoting the Gaussian kernels $\kappa_{1}$ and $\kappa_{2}$, the smoothed gradient value at $\theta_{t}$ should be as follows:
$g_{t}=\frac{\kappa_{1}(\theta_{t},\theta_{t})}{\kappa_{1}(\theta_{t},\theta_{t})+\kappa_{2}(\theta_{t},\theta_{t})}u_{t}+\frac{\kappa_{2}(\theta_{t},\theta_{t})}{\kappa_{1}(\theta_{t},\theta_{t})+\kappa_{2}(\theta_{t},\theta_{t})}\hat{g}_{t}.$ (67)
@@
Note that $u_{t}$ in VSGD is introduced as a learned summary of the noisy observations before $t$ and we only need to smooth on $u_{t}$ as a cheap approximation to smoothing on all of the historical data. Similar simplification technique is commonly seen in scalable Gaussian Process solutions, as discussed in Quinonero-Candela & Rasmussen (2005); Liu et al. (2020a). In these references, the learned summaries are called the "inducing points".
@@ A.3 Treating Samples Separately in a Mini-bach
Assuming the size of the mini-batch is $M$, denote the sample gradients $\{\hat{g}_{t}^{(i)}\}_{i=1:M}$. According to Hoffman et al. (2013), we would need to calculate Eq. 14 $M$ times and average them in Eq. 17 and Eq. 18. Specifically, we would need to replace Eq. 14, Eq. 17 and Eq. 18 with:
@@ B.3 Data Augmentations
B.4 Training Loss
@@
Figure 6: Top-1 Accuracy on the Imagenet dataset trained with ResNet-50.
Table 5: Final Average train loss, over three random seeds.
VSGD VSGD Adam AdamW SGD
(w/ L2) (w/o L2) (w/o L2) (w/ L2) (w/ mom)
CIFAR100
VGG16 0.0016 0.0022 0.0284 0.0157 0.0018
ConvMixer 0.0019 0.0022 0.0094 0.0094 0.0027
ResNeXt-18 0.0027 0.0015 0.0081 0.0088 0.0151
TinyImagenet-200
VGG19 0.8783 0.1900 1.2528 1.4722 0.1954
ConvMixer 0.8823 0.8513 0.7347 0.6882 0.7581
ResNeXt-18 1.3233 1.2680 1.4096 1.4089 1.8158